# Unifying Symbolic Music Arrangement: Track-Aware Reconstruction and Structured Tokenization

### v3

@@ A.2 Example
We name our tokenization scheme REMI-z for its distinctive “zig-zag” encoding pattern for musical notes within each bar, as illustrated in Figure 4(a). In this scheme, notes are encoded hierarchically: first grouped by tracks (instruments), then organized bar by bar. This track-first approach ensures notes from the same instrument remain clustered together, thereby enhancing the model’s ability to learn instrument-specific patterns. The resulting REMI-z sequence is demonstrated in Figure 4(b). In contrast, REMI+ [28] employs a column-wise encoding strategy, strictly ordering notes by their temporal positions. While this approach effectively captures global temporal relationships, it disperses notes from the same instrument throughout the sequence, potentially complicating instrument-specific pattern learning. We leverage this temporal ordering advantage of REMI+ in our content sequence (excluding instrument tokens), as shown in Figure 5.
@@ B.1 Key Normalization
Transposition-equivariance is a crucial property in symbolic music: a composition’s musicality remains unchanged under global pitch shifts (uniform pitch adjustment across all notes). However, uneven key distribution in datasets can lead to data sparsity during training, potentially causing models to perform inconsistently across different keys. Two approaches address this issue: (1) data augmentation through systematic semitone transpositions $0,\pm 1,\pm 2,\cdots$} [17, 39], or (2) normalizing all songs to a common key (e.g., C major and A minor) [15]. We adopt the latter approach, implementing a modified version of [15]’s method. Our implementation uses a key dictionary mapping from each of the 24 keys (12 major and 12 minor) to a 12-dimensional binary vectors, where 1s indicate scale notes. Key detection is performed by computing dot products between these vectors and a song’s pitch histogram, with the highest-scoring key determining the transposition needed to normalize to C major or A minor. Further details can be found in the code.
@@ B.2 Random Track Deletion
Concretely, given a content sequence with multiple instrument tracks, we randomly delete a subset of instruments before feeding the sequence to the model. The number of instruments to delete is sampled from a Poisson distribution with $\lambda=\max(\lfloor|\mathcal{I}|/4\rfloor,1)$, where $\mathcal{I}$ is the set of non-melodic instruments in the content sequence. The melodic track is estimated by the track with highest average pitch, and does not involves in the track deletion to ensure the melody’s content is preserved. Then, the deletion count is clipped to ensure at least one instrument remains. A corresponding number of instruments are uniformly sampled without replacement and all tokens associated with these instruments are removed from the input. Then the content sequence is calculated on this modified music sequence.
Importantly, the output target remains unchanged: it includes all instruments, including those that were deleted from the input. This setup requires the model to reconstruct the missing tracks based solely on the remaining musical context. By doing so, the model learns not only to replicate existing content, but also to infer what kind of notes would be appropriate for a given instrument in a specific musical context, even when such content is not explicitly present in the input.
@@ B.3 Instrument Quantization
While our tokenization scheme supports all MIDI program IDs, many instruments with different IDs share fundamental compositional properties, differing primarily in timbre (e.g., acoustic and electric pianos). To leverage these similarities and reduce instrument distribution sparsity in training data, we group similar instruments and assign them the lowest program ID within their group. For multiple tracks of the same instrument type, we merge their notes into a single track.
@@ B.4 Baseline Models
Tokenization Comparison
When comparing the effectiveness of REMI+ and the proposed REMI-z tokenization, the model structure used is the same as our arrangement model, but with a simpler training objective—the standard left-to-right next-token-prediction.
@@
For band arrangement, we use the Transformer-VAE model from [39] as our baseline, utilizing their official implementation without retraining. The model’s generation module was trained on Slakh2100 [19] (the same dataset as our fine-tuning) and its prior model leveraged the larger Lakh MIDI Dataset [22], which encompasses Slakh2100.
@@
Piano reductions can be categorized into two types: (1) piano accompaniments where the melody is delegated to a separate lead instrument, focusing solely on preserving harmony and texture (e.g., [38, 30]), which can be effectively handled using self-supervised methods; and (2) complete solo piano arrangements that additionally include the melody, requiring careful human arrangement and band-to-piano parallel datasets for supervised learning (e.g., [27, 26]). Due to the lack of open-source parallel datasets, we focus on the accompaniment arrangement task in this paper and use the term piano reduction interchangeably. However, our methodology is inherently flexible and can also handle the second type of reduction if parallel datasets become available. Additionally, exploring the potential for full solo piano reductions using unsupervised approaches remains a valuable direction for future research.
For comparison, we reimplemented the UNet baseline from [27] and trained it on Slakh2100 using our data preparation pipeline. Instead of adopting parallel data, we followed the same setting as our model, using the original composition as input and the piano track within the song as output. To ensure musical coverage, we selected only piano tracks that span more than 40% of the piece’s pitch range.
@@
For drum arrangement, we employ Composer’s Assistant v2 [17] as our baseline, specifically using v2.1.0 from their official repository [^2] https://github.com/m-malandro/composers-assistant-REAPER. We use the model without retraining since its original training data (Lakh MIDI Dataset [22]) encompasses our fine-tuning dataset (Slakh2100).
@@ B.5 Hyperparameter Settings
The model we adopted is a GPT-2 model comprising 12 Transformer decoder layers with a hidden size of 768. We detail the hyperparameter configurations used in our experiments below.
@@ C.1 Arrangement Evaluation
Finally, Voice WER (V-WER) evaluates the similarity of voice arrangements between two multitrack compositions. For each piece, we derive an ordered voice sequence by:
@@
The V-WER between output sequence $V^{out}$ and target sequence $V^{tgt}$ is:
$\text{V-WER}=\frac{S+D+I}{N}$ (12)
@@ D.2 Human Evaluation Details
Each sample group includes the original music to be arranged and the anonymized outputs from all compared models (plus the ground truth for the drum task). Each sample is 8–16 bars long, rendered to audio in Cubase AI 13 using the default soundfont and the original MIDI’s BPM. We collected 56, 73, and 77 evaluation groups for band, piano, and drum tasks respectively. A total of 26 participants joined the study, 5 of whom work in the music industry. Among all evaluators, 73.1% have over 10 years of experience in music composition or performance. The mean time spent per session was approximately 30 minutes. Figure 6 shows the sample survey interface and instructions.
@@
Evaluation Protocol. Participants first listened to the original music clip, followed by model outputs presented in random order. For each arranged clip, participants rated multiple aspects using a 5-point Likert scale, based on task-specific questions in the previous subsection. The evaluation groups were randomly distributed to participants to ensure unbiased feedback.
@@ Appendix E Probing Analysis of Pre-Training Impact
Knowledge probing techniques are used to explore what a model has learned within its hidden representations [24]. A widely adopted method is linear probing, where the pre-trained model is frozen and a simple linear classifier is trained on top of the hidden representations to predict specific properties, thereby revealing the extent to which particular knowledge is encoded in the model [1]. This technique is particularly useful in our context for evaluating what the model has internalized during pre-training.
@@ Appendix F Broader Impact
Beyond arrangement tasks, the proposed tokenization scheme may have broader implications for symbolic music modeling in general. By restructuring musical data into locally coherent, instrument-consistent sequences, it facilitates learning over meaningful musical units. This could improve both conditional and unconditional generation quality across tasks such as continuation, accompaniment generation, or reharmonization. Moreover, we anticipate that structured representations like REMI-z may benefit music understanding tasks, including symbolic transcription from audio, by facilitating modeling of instrument-wise playing styles when decoding. We encourage future work to explore these directions.
@@ Appendix G Limitations
Fixed creativity level in band arrangement. In the band arrangement task, we encourage creative rearrangements by randomly deleting tracks from the input content sequence during training. However, this leads to a somewhat fixed creativity level that does not adapt to the target instrumentation or musical context. In reality, different scenarios may require varying degrees of creativity—for example, arranging a simple piece for a full orchestra requires more elaboration than arranging it for a trio. Learning to adapt the level of creativity to the reinstrumentation context is an open challenge.

### v5

@@ A.2 Example
We name our tokenization scheme REMI-z for its distinctive “zig-zag” encoding pattern for musical notes within each bar, as illustrated in Figure 4(a). In this scheme, notes are encoded hierarchically: first grouped by tracks (instruments), then organized bar by bar. This track-first approach ensures notes from the same instrument remain clustered together, thereby enhancing the model’s ability to learn instrument-specific patterns. The resulting REMI-z sequence is demonstrated in Figure 4(b). In contrast, REMI+ [von2023figaro] employs a column-wise encoding strategy, strictly ordering notes by their temporal positions. While this approach effectively captures global temporal relationships, it disperses notes from the same instrument throughout the sequence, potentially complicating instrument-specific pattern learning. We leverage such temporal ordering in our content sequence (excluding instrument tokens), as shown in Figure 5.
@@ B.1 Key Normalization
Transposition-equivariance is a crucial property in symbolic music: a composition’s musicality remains unchanged under global pitch shifts (uniform pitch adjustment across all notes). However, uneven key distribution in datasets can lead to data sparsity during training, potentially causing models to perform inconsistently across different keys. Two approaches address this issue: (1) data augmentation through systematic semitone transpositions $0,\pm 1,\pm 2,\cdots$} [ComposersAssistant2, zhao2023accomontage], or (2) normalizing all songs to a common key (e.g., C major and A minor) [lu2023musecoco]. We adopt the latter approach, implementing a modified version of [lu2023musecoco]’s method. Our implementation uses a key dictionary mapping from each of the 24 keys (12 major and 12 minor) to a 12-dimensional binary vectors, where 1s indicate scale notes. Key detection is performed by computing dot products between these vectors and a song’s pitch histogram, with the highest-scoring key determining the transposition needed to normalize to C major or A minor. Further details can be found in the code.
@@ B.2 Random Track Deletion
Concretely, given a content sequence with multiple instrument tracks, we randomly delete all notes that belongs to a subset of instruments before calculating the content sequences and feeding to the model. The number of instruments to delete is sampled from a Poisson distribution with $\lambda=\max(\lfloor|\mathcal{I}|/4\rfloor,1)$, where $\mathcal{I}$ is the set of non-melodic instruments in the content sequence. The melodic track is estimated by the track with highest average pitch, and does not involves in the track deletion to ensure the melody’s content is preserved. Then, the deletion count is clipped to ensure at least one instrument remains. A corresponding number of instruments are uniformly sampled without replacement and all tokens associated with these instruments are removed from the input. Then the content sequence is calculated on this modified music sequence.
Importantly, the target sequence remains unchanged: it includes all instruments and notes removed from the input. This setup requires the model to reconstruct missing tracks based solely on the remaining musical context. In doing so, the model learns not only to replicate observed content but also to infer plausible notes for desired instruments in a given context. This can be treated as a music specific span infilling denoising objective for sequence-to-sequence learning [lewis2019bart], but we don’t explicitly tell the model whether an input is masked and the location of masked spans.
@@ B.3 Instrument Quantization
While our tokenization scheme supports all MIDI program IDs, many instruments with different IDs share fundamental compositional properties, differing primarily in timbre (e.g., acoustic and electric pianos). To leverage these similarities and reduce instrument distribution sparsity in training data, during fine-tuning, we group similar instruments and assign them the lowest program ID within their group. The instrument grouping rules follows that of [manilow2019cutting], leading to 34 different program IDs in total. For multiple tracks of the same instrument type, we merge their notes into a single track.
@@ B.4 Baseline Models
For band arrangement, we use the Transformer-VAE model from [zhao2023accomontage] as our baseline. For human evaluation, we utilized their official implementation without retraining. The model’s generation module was trained on Slakh2100 [manilow2019cutting] (the same dataset as our fine-tuning), and its prior model leveraged the larger Lakh MIDI Dataset [Raffel2016], which encompasses Slakh2100. For objective evaluation, we retrained another version of the model on Slakh2100 that did not receive hints from the style prior model for a fair comparison.
@@
Piano reductions can be categorized into two types: (1) piano accompaniments where the melody is delegated to a separate lead instrument, focusing solely on preserving harmony and texture (e.g., [zhao2023q, wang2020pop909]), which can be effectively handled using self-supervised methods; and (2) complete solo piano arrangements that additionally include the melody, requiring careful human arrangement and band-to-piano parallel datasets for supervised learning (e.g., [terao2023neural, terao2022difficulty]). Due to the lack of open-source parallel datasets, we focus on the accompaniment arrangement task in this paper and use the term piano reduction interchangeably. However, our methodology is inherently flexible and can also handle the second type of reduction if parallel datasets become available. Additionally, exploring the potential for full solo piano reductions using unsupervised approaches remains a valuable direction for future research.
For comparison, we reimplemented the UNet baseline from [terao2023neural] and trained it on Slakh2100 using our data preparation pipeline. Instead of adopting parallel data, we followed the same setting as our model, using the original composition as input and the piano track within the song as output. To ensure musical coverage, we selected only piano tracks that span more than 40% of the piece’s pitch range.
@@
For drum arrangement, we employ Composer’s Assistant v2 [ComposersAssistant2] as our baseline, specifically using v2.1.0 from their official repository [^2] https://github.com/m-malandro/composers-assistant-REAPER. We use the model without retraining since its original training data (Lakh MIDI Dataset [Raffel2016]) encompasses our fine-tuning dataset (Slakh2100).
Tokenization Comparison
When comparing the effectiveness of REMI+ and the proposed REMI-z tokenization in §3.2, the model structure used is the same as our arrangement model, but with a simpler training objective—the standard left-to-right next-token-prediction. We did not use time signature tokens, tempo tokens, and velocity tokens in REMI+ for a fair comparison.
@@ B.5 Hyperparameter Settings
The model we adopted is a GPT-2 model comprising 12 Transformer decoder layers with a hidden size of 768. Our setup follows GPT’s 12-layer, 768-dim embedding, 3072-dim inner states [radford2018improving], and our experimental design is also in the same style: a single pretrained model and multiple fine-tuning tasks to show the efficacy of the pretrain–finetune paradigm. We used a slightly higher number of attention heads (16 instead of 12) based on the intuition that the interactions between different music tokens are more critical for musical quality than the value of individual note attributes. We detail the hyperparameter configurations used in our experiments below.
@@ C.1 Arrangement Evaluation
Finally, Voice Error Rate (VER) evaluates the similarity of voice arrangements between two multitrack compositions. For each piece, we derive an ordered voice sequence by:
@@
The VER between output sequence $V^{out}$ and target sequence $V^{tgt}$ is:
$\text{VER}=\frac{S+D+I}{N}$ (12)
@@
Among all the selected objective metrics, the Notei F1 is the most comprehensive and important one. To achieve a perfect Notei F1 score, the model need to perform perfectly in all evaluated aspects: instruments, voice, melody, note reconstruction and allocation.
@@ D.2 Human Evaluation Details
Each sample group includes the original music to be arranged and the anonymized outputs from all compared models (plus the ground truth for the drum task). Each sample is 8–16 bars long, rendered to audio in Cubase AI 13 using the default soundfont and the original MIDI’s BPM. We collected 56, 73, and 77 evaluation groups for band, piano, and drum tasks respectively. The mean time spent per session was approximately 30 minutes. Figure 6 shows the sample survey interface and instructions.
@@
Participant Background A total of 26 participants joined the study, 5 of whom work in the music industry. Among all evaluators, 73.1% have over 10 years of experience in music composition or performance. We conducted t-tests comparing these raters with those having $\leq$10 years of experience (using piano reduction ratings on our model), and found no significant differences in the means between the two groups ($p>0.2$ across all metrics). This suggests that musical background did not systematically bias the results.
Evaluation Protocol. Participants first listened to the original music clip, followed by model outputs presented in random order. For each arranged clip, participants rated multiple aspects using a 5-point Likert scale, based on task-specific questions in the previous subsection. The evaluation groups were randomly distributed to participants to ensure unbiased feedback. Each rater evaluated a subset of groups, where one group = one song (input) with multiple model outputs. On average, raters evaluated 2.15 groups for band, 2.80 for piano, and 2.96 for drum — corresponding to 8.6, 14.0, and 11.8 samples respectively.
@@
D.3 Discussion on Evaluation Scheme
Our experiments primarily adopt a 5-point scale for subjective evaluation, which is the most commonly used protocol in prior music arrangement works [zhao2023accomontage, wang2022audio, yi2022accomontage2]. However, we observe a growing trend toward using A/B testing to evaluate overall quality in the broader music generation literature [wang2025notagen, qu2024mupt, wu2024melodyt5]. We encourage future research to consider combining both approaches to obtain more comprehensive and convincing evaluation results.
@@ Appendix E Probing Analysis of Pre-Training Impact
Knowledge probing techniques are used to explore what a model has learned within its hidden representations [rogers2020primer]. A widely adopted method is linear probing, where the pre-trained model is frozen and a simple linear classifier is trained on top of the hidden representations to predict specific properties, thereby revealing the extent to which particular knowledge is encoded in the model [alain2016understanding]. This technique is particularly useful in our context for evaluating what the model has internalized during pre-training.
@@ Appendix F Broader Impact
Many other conditional generation tasks could similarly benefit from this training objective design. For instance, one can construct a conditional sequence derived from a partial or aspect-wise decomposition of the target, concatenate it with the original target tokens, and then perform segment-level generation using a pretrained model with historical context-awareness to fine-tune existing generative models. Examples include bar-level infilling (e.g., removing and rewriting one bar), melody generation conditioned on chords, variation generation from a given melody, counter-melody generation, or harmonizing a melody with chords. All of these tasks share the same music-conditioned generation pattern and may benefit from a strong generative pretrained model.
Beyond arrangement tasks, the proposed tokenization scheme may have broader implications for symbolic music modeling in general. By restructuring musical data into continuous track sequences, it facilitates learning over reusable musical phrases. This may improve both conditional and unconditional generation quality across tasks such as continuation, accompaniment generation, and reharmonization. Moreover, we anticipate that structured representations like REMI-z could benefit music understanding tasks, including symbolic transcription from audio, by facilitating the modeling of instrument-wise playing styles during decoding. Additionally, one may consider applying sequence compression techniques—such as Byte Pair Encoding (BPE) or variational autoencoders (VAE)—on top of this base tokenization to enhance compactness, which could further support long-range modeling essential for full-song generation. We encourage future work to explore these directions.
Beyond arrangement tasks, the proposed tokenization scheme may have broader implications for symbolic music modeling in general. By restructuring musical data into locally coherent, instrument-consistent sequences, it facilitates learning over meaningful musical units. This could improve both conditional and unconditional generation quality across tasks such as continuation, accompaniment generation, or reharmonization. Moreover, we anticipate that structured representations like REMI-z may benefit music understanding tasks, including symbolic transcription from audio, by facilitating modeling of instrument-wise playing styles when decoding. Also, you can consider combine sequence compression techniques, maybe BPE, maybe VAE, on top of that, to further enhance the compactness, which may facilitate long-range modeling that are important to full-song generation. We encourage future work to explore these directions.
@@ Appendix G Limitations
Handling of Time Signature and Tempo. Our current model does not explicitly model time signature or tempo tokens during fine-tuning. As a result, it is limited to handling music with a fixed 4/4 time signature and stable tempo. This design choice follows common practice in prior arrangement research [zhao2023accomontage, wang2022audio, terao2023neural], where similar assumptions are made to simplify modeling. Empirically, approximately 95% of the fine-tuning data is in 4/4 time (see Appendix A), and stable tempo is generally the norm in modern music [dannenberg2011characterizing]. Nevertheless, extending the framework to handle non-4/4 time signatures and dynamic tempo variations remains an important direction for future work.